Melly bought a single storey house in Taman Nasional for RM230,000. She

made a 10% down payment and financed the balance through a bank loan from

Maybank for 30 years.

a) Find her monthly payment R if the interest rate was 7.5% compounded

monthly. How much was the total interest charged?

[6 Marks]

b) Due to COVID-19 pandemic, Melly missed the first three payments. How

much should be paid on the fourth month if she wanted to settle the

outstanding arrears?

[3 Marks]

c) Melly decided to settle the loan in full immediately after paying for 10 years.

How much is the amount that needs to be paid? 


1
Expert's answer
2021-02-25T01:52:21-0500

n= 30 x 12 =360 years

i = 0.075/12

0.00625%

pv = 230000-(10% of 230000)

= 207000

PV=P(1-(1-i)-n/i


207000=p(1-(1-0.00625)-power 360/0.00625

207000=-1368.64 p

151.24

b) 151.24 x 4

604.96

c) 10x12x151.24 =18148.8


207000-18148.8

=188851.2